OpenStudy (anonymous):

A ball thrown horizontally at 20 m/s travels a horizontal distance of 46 m before hitting the ground. From what height was the ball thrown?

OpenStudy (anonymous):

 1st you have to find how the ball to travel the horizontal component. distance/velocity=time 46/20=2.3 2nd you have to find vertical component of the ball falling Velocity=time*acceleration 2.3*9.81=22.563 3rd now you have velocity and how long it fell. so now you plug those number into the formula distance=velocity*time distance=22.563*2.3 hope this helps. Im in Physics C so this is the easy stuff lol. prepare yourself for the rest of the year...